Adult tickets for a play cost $6 and child tickets cost $1. If there were 32 people at a performance and the theater collected $117 from ticket sales, how many children attended the play?
Answer by scott8148(6628)

let x=children, so 32-x=adults

1(x)+6(32-x)=117 ___ x+192-6x=117 ___ -5x=-75